OpenStudy (anonymous):

[4.07] Choose the slope-intercept equation of the line that passes through the point (6, 2) and is perpendicular to y = 3x + 6.

OpenStudy (anonymous):

@eric_d

OpenStudy (eric_d):

Y=3x+6 m=3 A line perpendicular to it will have slope -1/3

OpenStudy (anonymous):

y = x y = -x + 4 y = 3x + 20 y = -3x - 16 theseare the choices @eric_d

OpenStudy (eric_d):

(y - 2) / (x - 6) = -1/3

OpenStudy (eric_d):

y - 2 = (-1/3)(x - 6) y - 2 = (-1/3)x + 2 y = (-1/3)x + 4

OpenStudy (eric_d):



OpenStudy (anonymous):

y=1/3x y=-1/3x +4 y= 3x+20 y=-3x-16 @eric_d

OpenStudy (eric_d):

Option 2

OpenStudy (eric_d):

I had shown you the full working

OpenStudy (eric_d):

Get me @kylieedeckahh

OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-2, 2) and is parallel to y = 4x + 7. y = 4x + 10 y= 1/4x+5/2 y=-4x-6 y=-1/4x+3/2 @eric_d

OpenStudy (eric_d):

When parallel.. both equations have the same gradient

OpenStudy (anonymous):

what does that mean @eric_d

OpenStudy (eric_d):

@nirmalnema Hope u can explain.. thanks in advance